OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3) Solve the system of equations. 2x - 5y = 3 x - 3y = 1 (4, 1) (5, 2) (7, 2) (9, 3) please help me and explain possibly?

OpenStudy (alexandervonhumboldt2):

sok so first y=1+2x and y=-1+4x. thus 1+2x=-1+4x now just solve for x: 2=2x. x=1. now substitue x inot one of the equations and solve for y: -2*1+y=1 y=3 x=1 y=3

OpenStudy (alexandervonhumboldt2):

to solve the second one first subtartc second equatino from first: 2x-5y-x+3y=3-1 x-2y=2 x=2+2y.

OpenStudy (alexandervonhumboldt2):

then solve for x=2+2y and x=1+3y

OpenStudy (alexandervonhumboldt2):

1+3y=2+2y y=1 now substitue y inot one of the equations and solve for x. as the multiply choice question there is only one choice A with y=1